\begin{thebibliography}{10}

\bibitem{981648}
K.K. Aggarwal, Y.~Singh, and J.K. Chhabra.
\newblock An integrated measure of software maintainability.
\newblock In {\em Reliability and Maintainability Symposium, 2002. Proceedings.
  Annual}, pages 235 --241, 2002.

\bibitem{antoniol2002recovering}
Giuliano Antoniol, Gerardo Canfora, Gerardo Casazza, Andrea De~Lucia, and
  Ettore Merlo.
\newblock Recovering traceability links between code and documentation.
\newblock {\em IEEE transactions on software engineering}, 28(10):970--983,
  2002.

\bibitem{binkley2011improving}
Dave Binkley, Matthew Hearn, and Dawn Lawrie.
\newblock Improving identifier informativeness using part of speech
  information.
\newblock In {\em Proceedings of the 8th Working Conference on Mining Software
  Repositories}, pages 203--206. ACM, 2011.

\bibitem{budanitsky2006evaluating}
Alexander Budanitsky and Graeme Hirst.
\newblock Evaluating wordnet-based measures of lexical semantic relatedness.
\newblock {\em Computational Linguistics}, 32(1):13--47, 2006.

\bibitem{7884638}
P.~Chatterjee, M.~A. Nishi, K.~Damevski, V.~Augustine, L.~Pollock, and N.~A.
  Kraft.
\newblock What information about code snippets is available in different
  software-related documents? an exploratory study.
\newblock In {\em 2017 IEEE 24th International Conference on Software Analysis,
  Evolution and Reengineering (SANER)}, pages 382--386, Feb 2017.

\bibitem{Chatterjee:2017}
Preetha Chatterjee, Benjamin Gause, Hunter Hedinger, and Lori Pollock.
\newblock Extracting code segments and their descriptions from research
  articles.
\newblock In {\em Proceedings of the 14th International Conference on Mining
  Software Repositories}, MSR '17, pages 91--101. IEEE Press, 2017.

\bibitem{Deimel:1985}
Lionel~E. Deimel, Jr.
\newblock The uses of program reading.
\newblock {\em SIGCSE Bull.}, 17(2):5--14, June 1985.

\bibitem{enslen2009mining}
Eric Enslen, Emily Hill, Lori Pollock, and K~Vijay-Shanker.
\newblock Mining source code to automatically split identifiers for software
  analysis.
\newblock In {\em Mining Software Repositories, 2009. MSR'09. 6th IEEE
  International Working Conference on}, pages 71--80. IEEE, 2009.

\bibitem{erlikh00}
Len Erlikh.
\newblock Leveraging legacy system dollars for e-business.
\newblock {\em IT Professional}, 2(3):17--23, 2000.

\bibitem{falleri2010automatic}
J-R Falleri, Marianne Huchard, Mathieu Lafourcade, Clementine Nebut, Violaine
  Prince, and Michel Dao.
\newblock Automatic extraction of a wordnet-like identifier network from
  software.
\newblock In {\em Proceedings of the 18th International Conference on Program
  Comprehension (ICPC)}, pages 4--13. IEEE, 2010.

\bibitem{goldberg-1987}
A.~Goldberg.
\newblock {Programmer as Reader}.
\newblock {\em IEEE Softw.}, 4(5):62--70, 1987.

\bibitem{Haiduc}
Sonia Haiduc, Jairo Aponte, and Andrian Marcus.
\newblock Supporting program comprehension with source code summarization.
\newblock In {\em Proceedings of the 32nd ACM/IEEE International Conference on
  Software Engineering - Volume 2}, ICSE '10, pages 223--226, New York, NY,
  USA, 2010. ACM.

\bibitem{haiduc2013automatic}
Sonia Haiduc, Gabriele Bavota, Andrian Marcus, Rocco Oliveto, Andrea De~Lucia,
  and Tim Menzies.
\newblock Automatic query reformulations for text retrieval in software
  engineering.
\newblock In {\em Proceedings of the 2013 International Conference on Software
  Engineering}, pages 842--851. IEEE Press, 2013.

\bibitem{hill2009automatically}
Emily Hill, Lori Pollock, and K~Vijay-Shanker.
\newblock Automatically capturing source code context of nl-queries for
  software maintenance and reuse.
\newblock In {\em Proceedings of the 31st International Conference on Software
  Engineering}, pages 232--242. IEEE Computer Society, 2009.

\bibitem{Jackson200840}
Stoney Jackson, Premkumar Devanbu, and Kwan-Liu Ma.
\newblock Stable, flexible, peephole pretty-printing.
\newblock {\em Science of Computer Programming}, 72(1-2):40 -- 51, 2008.
\newblock Special Issue on Second issue of experimental software and toolkits
  (EST).

\bibitem{kuhn07}
Adrian Kuhn, St\'{e}phane Ducasse, and Tudor G\'{\i}rba.
\newblock Semantic clustering: Identifying topics in source code.
\newblock {\em Inf. Softw. Technol.}, 49(3):230--243, 2007.

\bibitem{lawrie2011expanding}
Dawn Lawrie and Dave Binkley.
\newblock Expanding identifiers to normalize source code vocabulary.
\newblock In {\em Proceedings of the 27th IEEE International Conference on
  Software Maintenance (ICSM)}, pages 113--122. IEEE, 2011.

\bibitem{lientz78}
B.~P. Lientz, E.~B. Swanson, and G.~E. Tompkins.
\newblock Characteristics of application software maintenance.
\newblock {\em Commun. ACM}, 21(6):466--471, 1978.

\bibitem{maarek1991information}
Yo{\"e}lle~S Maarek, Daniel~M Berry, and Gail~E Kaiser.
\newblock An information retrieval approach for automatically constructing
  software libraries.
\newblock {\em IEEE Transactions on software Engineering}, 17(8):800--813,
  1991.

\bibitem{madani2010recognizing}
Nioosha Madani, Latifa Guerrouj, Massimiliano Di~Penta, Yann-Gael Gueheneuc,
  and Giuliano Antoniol.
\newblock Recognizing words from source code identifiers using speech
  recognition techniques.
\newblock In {\em Proceedings of the 14th European Conference on Software
  Maintenance and Reengineering (CSMR)}, pages 68--77. IEEE, 2010.

\bibitem{manning1999foundations}
Christopher~D Manning and Hinrich Sch{\"u}tze.
\newblock {\em Foundations of statistical natural language processing}, volume
  999.
\newblock MIT Press, 1999.

\bibitem{marcus2003recovering}
Andrian Marcus and Jonathan~I Maletic.
\newblock Recovering documentation-to-source-code traceability links using
  latent semantic indexing.
\newblock In {\em Software Engineering, 2003. Proceedings. 25th International
  Conference on}, pages 125--135. IEEE, 2003.

\bibitem{marcus2004information}
Andrian Marcus, Andrey Sergeyev, Vaclav Rajlich, and Jonathan~I Maletic.
\newblock An information retrieval approach to concept location in source code.
\newblock In {\em Reverse Engineering, 2004. Proceedings. 11th Working
  Conference on}, pages 214--223. IEEE, 2004.

\bibitem{martin2000speech}
James~H Martin and Daniel Jurafsky.
\newblock Speech and language processing.
\newblock {\em International Edition}, 710, 2000.

\bibitem{Maskeri2008}
Girish Maskeri, Santonu Sarkar, and Kenneth Heafield.
\newblock Mining business topics in source code using latent dirichlet
  allocation.
\newblock In {\em Proceedings of the 1st India Software Engineering Conference
  (ISEC)}, pages 113--120. ACM, 2008.

\bibitem{6613830}
L.~Moreno, J.~Aponte, G.~Sridhara, A.~Marcus, L.~Pollock, and K.~Vijay-Shanker.
\newblock Automatic generation of natural language summaries for java classes.
\newblock In {\em Program Comprehension (ICPC), 2013 IEEE 21st International
  Conference on}, pages 23--32, 2013.

\bibitem{murphy-ecoop-05}
Gail Murphy, Mik Kersten, Martin Robillard, and Davor Cubranic.
\newblock The emergent structure of development tasks.
\newblock In {\em ECOOP 2005 --- Object-Oriented Programming, 19th European
  Conf}, Glasgow, Scotland, July~27--29, 2005.

\bibitem{Ohba2005}
Masaru Ohba and Katsuhiko Gondow.
\newblock Toward mining "concept keywords" from identifiers in large software
  projects.
\newblock In {\em Proceedings of the 2005 International Workshop on Mining
  Software Repositories (MSR)}, pages 1--5. ACM, 2005.

\bibitem{pedersen2005maximizing}
Ted Pedersen, Satanjeev Banerjee, and Siddharth Patwardhan.
\newblock Maximizing semantic relatedness to perform word sense disambiguation.
\newblock {\em University of Minnesota supercomputing institute research report
  UMSI}, 25:2005, 2005.

\bibitem{peng2013improving}
Xin Peng, Zhenchang Xing, Xi~Tan, Yijun Yu, and Wenyun Zhao.
\newblock Improving feature location using structural similarity and iterative
  graph mapping.
\newblock {\em Journal of Systems and Software}, 86(3):664--676, 2013.

\bibitem{petrenko2013concept}
Maksym Petrenko and V{\'a}Clav Rajlich.
\newblock Concept location using program dependencies and information retrieval
  (depir).
\newblock {\em Information and Software Technology}, 55(4):651--659, 2013.

\bibitem{Rajlich:2004}
Vaclav Rajlich and Prashant Gosavi.
\newblock Incremental change in object-oriented programming.
\newblock {\em IEEE Softw.}, 21(4):62--69, July 2004.

\bibitem{Raymond}
Darrell~R. Raymond.
\newblock Reading source code.
\newblock In {\em Proceedings of the 1991 conference of the Centre for Advanced
  Studies on Collaborative research}, CASCON '91, pages 3--16. IBM Press, 1991.

\bibitem{readingcode3}
Spencer Rugaber.
\newblock The use of domain knowledge in program understanding.
\newblock {\em Annals of Software Engineering}, 9(1-2):143--192, 2000.

\bibitem{schmid2013probabilistic}
Helmut Schmid.
\newblock Probabilistic part-ofispeech tagging using decision trees.
\newblock In {\em New methods in language processing}, page 154. Routledge,
  2013.

\bibitem{shepherd07}
David Shepherd, Zachary~P. Fry, Emily Hill, Lori Pollock, and K.~Vijay-Shanker.
\newblock Using {N}atural {L}anguage {P}rogram {A}nalysis to {L}ocate and
  {U}nderstand {A}ction-oriented {C}oncerns.
\newblock In {\em AOSD '07: 6th Intl Conf on Aspect-oriented Software
  Development}, New York, NY, USA, 2007. ACM Press.

\bibitem{sridhara2010towards}
Giriprasad Sridhara, Emily Hill, Divya Muppaneni, Lori Pollock, and
  K~Vijay-Shanker.
\newblock Towards automatically generating summary comments for java methods.
\newblock In {\em Proceedings of the IEEE/ACM international conference on
  Automated software engineering}, pages 43--52. ACM, 2010.

\bibitem{giri08}
Giriprasad Sridhara, Emily Hill, Lori Pollock, and K~Vijay-Shanker.
\newblock Identifying {W}ord {R}elations in {S}oftware: {A} {C}omparative
  {S}tudy of {S}emantic {S}imilarity {T}ools.
\newblock In {\em 16th IEEE Intl Conf on Program Comprehension, 2008}, 2008.

\bibitem{giri11}
Giriprasad Sridhara, Lori Pollock, and K.~Vijay-Shanker.
\newblock Automatically detecting and describing high level actions within
  methods.
\newblock In {\em Proceedings of the International Conference on Software
  Engineering (ICSE)}, 2011.

\bibitem{sridhara2011automatically}
Giriprasad Sridhara, Lori Pollock, and K~Vijay-Shanker.
\newblock Automatically detecting and describing high level actions within
  methods.
\newblock In {\em Software Engineering (ICSE), 2011 33rd International
  Conference on}, pages 101--110. IEEE, 2011.

\bibitem{5970165}
Giriprasad Sridhara, Lori Pollock, and K~Vijay-Shanker.
\newblock Generating parameter comments and integrating with method summaries.
\newblock In {\em Program Comprehension (ICPC), 2011 IEEE 19th International
  Conference on}, pages 71--80. IEEE, 2011.

\bibitem{prog-comp-survey-05}
M.-A. Storey.
\newblock Theories, methods and tools in program comprehension: Past, present
  and future.
\newblock {\em 13th Intl Workshop on Program Comprehension, 2005}.

\bibitem{tian2014automated}
Yuan Tian, David Lo, and Julia Lawall.
\newblock Automated construction of a software-specific word similarity
  database.
\newblock In {\em Proceedings of the Conference on Software Maintenance,
  Reengineering and Reverse Engineering (CSMR-WCRE)}, pages 44--53. IEEE, 2014.

\bibitem{fifteen-years-of-pc}
Scott Tilley.
\newblock 15 years of program comprehension.
\newblock In {\em ICPC '07: Proceedings of the 15th IEEE International
  Conference on Program Comprehension}, pages 279--280, Washington, DC, USA,
  2007. IEEE Computer Society.

\bibitem{wang2013improving}
Jinshui Wang, Xin Peng, Zhenchang Xing, and Wenyun Zhao.
\newblock Improving feature location practice with multi-faceted interactive
  exploration.
\newblock In {\em Proceedings of the 2013 International Conference on Software
  Engineering}, pages 762--771. IEEE Press, 2013.

\bibitem{wang-orau}
X.~Wang, L.~Pollock, and K.~Vijay-Shanker.
\newblock Automatically generating natural language descriptions for
  object-related statement sequences.
\newblock In {\em 2017 IEEE 24th International Conference on Software Analysis,
  Evolution and Reengineering (SANER)}, pages 205--216, Feb 2017.

\bibitem{wang-thesis}
Xiaoran Wang.
\newblock {\em Exploring Action Unit Granularity of Source Code for Supporting
  Software Maintenance}.
\newblock PhD thesis, University of Delaware, 11 2017.

\bibitem{wang2008categorization}
Xiaoran Wang and Coskun Bayrak.
\newblock Categorization of users using unlabeled query logs.
\newblock In {\em DMIN}, pages 674--679, 2008.

\bibitem{wang2011automatic}
Xiaoran Wang, Lori Pollock, and K~Vijay-Shanker.
\newblock Automatic segmentation of method code into meaningful blocks to
  improve readability.
\newblock In {\em Reverse Engineering (WCRE), 2011 18th Working Conference on},
  pages 35--44. IEEE, 2011.

\bibitem{wang2013automatic}
Xiaoran Wang, Lori Pollock, and K~Vijay-Shanker.
\newblock Automatic segmentation of method code into meaningful blocks: Design
  and evaluation.
\newblock {\em Journal of Software: Evolution and Process}, 2013.

\bibitem{wang-2017-orau}
Xiaoran Wang, Lori Pollock, and K~Vijay-Shanker.
\newblock Automatically generating natural language descriptions for
  object-related statement sequences.
\newblock In {\em Proceedings of the 24th International Conference on Software
  Analysis, Evolution and Reengineering (SANER)}, pages 205--216, Feb 2017.

\bibitem{6693113}
E.~Wong, Jinqiu Yang, and Lin Tan.
\newblock Autocomment: Mining question and answer sites for automatic comment
  generation.
\newblock In {\em Automated Software Engineering (ASE), 2013 IEEE/ACM 28th
  International Conference on}, pages 562--567, Nov 2013.

\bibitem{wu2009learning}
Lei Wu, Linjun Yang, Nenghai Yu, and Xian-Sheng Hua.
\newblock Learning to tag.
\newblock In {\em Proceedings of the 18th international conference on World
  wide web}, pages 361--370. ACM, 2009.

\bibitem{Ying:2013}
Annie T.~T. Ying and Martin~P. Robillard.
\newblock Code fragment summarization.
\newblock In {\em Proceedings of the 2013 9th Joint Meeting on Foundations of
  Software Engineering}, ESEC/FSE 2013, pages 655--658, New York, NY, USA,
  2013. ACM.

\bibitem{zhao2006sniafl}
Wei Zhao, Lu~Zhang, Yin Liu, Jiasu Sun, and Fuqing Yang.
\newblock Sniafl: Towards a static noninteractive approach to feature location.
\newblock {\em ACM Transactions on Software Engineering and Methodology
  (TOSEM)}, 15(2):195--226, 2006.

\end{thebibliography}

\documentclass[conference]{IEEEtran}

\usepackage{url}
\usepackage{graphicx}
\usepackage{multirow}   
\usepackage{enumitem}
\usepackage{xspace}  
\usepackage{verbatim}   
\usepackage{algpseudocode}
\usepackage{lipsum}
\usepackage{float}
\usepackage{graphicx}
\usepackage{flushend}
\usepackage{caption}
\usepackage{changepage}
\usepackage{scrextend}
\usepackage[square,numbers]{natbib}

\begin{document}
\title{Comment Generation for Source Code: Survey}

\author{\IEEEauthorblockN{Xiaoran Wang and Benwen Zhang}
\IEEEauthorblockA{
xiaoranlr@gmail.com, benwen@udel.edu}
}

\maketitle

\begin{abstract}
Research has shown that most effort of today's software development is maintenance and evolution. Developers often use integrated development environments, debuggers, and tools for code search, testing, and program understanding to reduce the tedious tasks. One way to make software development more efficient is to make the program more readable. There have been many approaches proposed and developed for this purpose. Among these approaches, comment generation for source code is gaining more and more attention and has become a popular research area. In this paper, the state of art in comment generation research area are summarized and the challenges and future opportunities are discussed.

\end{abstract}

\begin{IEEEkeywords}
Mining software repositories, comment generation, natural langauge processing
\end{IEEEkeywords}

\section{Introduction}

Research has shown that more than 60\% of software engineering resources are spent on maintenance~\cite{erlikh00}. Software maintenance is the process of modifying a software system after delivery to fix bugs, improve performance, or adapt to a changing environment~\cite{lientz78}. Software maintenance requires code comprehension~\cite{981648}, as reading and understanding source code is the prerequisites of any modification. Program comprehension is time-consuming and cost most of developers' time~\cite{goldberg-1987,murphy-ecoop-05,Haiduc,Deimel:1985,Raymond,readingcode3}.  Developers often use integrated development environments, debuggers, and tools for code search, testing, and program understanding to reduce the tedious tasks. 
  
To reduce the effort of program comprehension and software maintenance, many tools have been developed to support the tedious and error-prone tasks~\cite{prog-comp-survey-05, fifteen-years-of-pc}. Formatting source code naturally helps program comprehension and there has been some research toward this direction~\cite{wang2011automatic,wang2013automatic,Jackson200840,wang-thesis}. More prevalently, researches have been conducted to generate comments to help program comprehension~\cite{sridhara2010towards,giri11,wang-2017-orau}. Comment generation for source code is gaining more and more attention and has become a popular research area.
In this survey, the state of art in the comment generation research area is summarized.

\section{State of the Art}

The researches of comment generation for source code can be categorized to three main classes based on the main technique utilized: 1) informaiton retrieval, 2) program structure information and 3) software artifacts beyond source code. This section first describes the work under the three categories and then summarize the fundamental natural langauge processing research in software engineering.

\subsection{Information Retrieval}

Researchers have integrated natural language processing techniques to analyze the natural language clues in the source code in recent years~\cite{lawrie2011expanding,
enslen2009mining,madani2010recognizing}. 
Information retrieval(IR) techniques use words in documents to determine the similarity between documents and queries. IR calculates a similarity score between a query and documents and the results of a query according to relevance are ranked based on the scores~\cite{shepherd07,wang2008categorization}. 

IR has become a standard technique to address the problem in natural language domain. Since its great promise in natural language, many researchers have applied IR to locate concepts~\cite{Rajlich:2004,zhao2006sniafl,haiduc2013automatic,hill2009automatically,marcus2004information,petrenko2013concept,peng2013improving,wang2013improving} and reconstruct documentation traceability links in source code~\cite{marcus2003recovering,antoniol2002recovering,Maskeri2008,Ohba2005}. These work use different variations of IR techniques to link the query with source code or build links between different software artifacts. The concepts located and links built can be served as the comment for the linked source code. 
IR-based concern location tools treat source code as a ``bag of words'' and focus on the individual words in the code. Research that uses IR does not generate natural langauge phrases or sentences. Instead, the generated comments are a bag of words. 

There have been many research towards extracting topic words or verb phrases from source code~\cite{Maskeri2008,Ohba2005,hill2009automatically}. They want to identify code fragments that are related to a given action or topic. Other research tries to cluster program elements that share similar phrases~\cite{kuhn07}.  These methods rely solely on the linguistic information to determine the topic of the code segment, which is often not adequate for describing the source code. 

Because IR approaches treat source code as a bag of words, the structure information is missing. More importantly, towards this research direction, the comment generated are not real ``comment''. At least, not like the comment developers put in the source code. 

\subsection{Program Structure Information}

In 2010, Sridhara et~al.~\cite{sridhara2010towards} first introduce an approach to generate summary comments for Java methods using program structure information. The basic idea is selecting the important statements from a method and then translate the statements into natural language phrases. They first build data flow graph of the method and then analyze the data flow to identify the important statements.

Later, Sridhara et~al.~\cite{sridhara2011automatically} introduced high level action in 2011. A high level action is defined as a high level step of what the code does. They identify a known set of high level actions based on a set of predefined templates that frequently occur in the source code. While the approach identifies code fragments that implement high level algorithmic steps, 
 the technique is limited in the kinds of high level action it is able to identify. Only 24\% of switch blocks, 40\% of if-else blocks, and  15\% of iterator loops implemented one of the templates. Later, Wang et al.~\cite{wang2011automatic} propose an automatic approach to cover loops without manually developing rules. They focus on identifying high level actions that are implemented by loop structures. While the work of Wang et al. advances the techniques, their technique is limited to loops only. After that, the same authors developed an approach to cover the more prevalent code fragments in source code - data flow chain~\cite{wang-orau}.
 
Complementary to the work of Sridhara et~al.~\cite{sridhara2010towards}, the same authors also presented a technique to generate comments for parameters and integrated those descriptions with method summaries~\cite{5970165}. Beyond method level, Moreno et al. developed an approach that generates summaries for Java classes~\cite{6613830}.

\subsection{Software Artifacts Beyond Source Code} 
The work that falls into ``program structure information'' category only uses the source code as the learning source to develop rules or build models. There have been some researches using forums to get the natural language clues, as developers often discuss the problem online in the site like Stack Overflow~\cite{7884638,Chatterjee:2017}..

Wong et al. used question and answer sites for automatic comment generation~\cite{6693113}. They extracted code-description pairs from the question title and text and used code clone detection to find source code snippets that are almost identical to the code of the code-description mappings. The code fragment in the pairs are often code copied from developers' projects and posted by developers. The basic idea is identifying comments for code fragments and reusing the comments for other projects. However, this technique only works for a small fraction of code fragments. The approach is scalable to millions of projects, but would not generate many comments for an individual project.

Ying and Robillard presented a supervised machine learning approach that classifies whether a line in a code fragment should be in a method summary~\cite{Ying:2013}. However, the generated summary is not the natural language but a set of statements.

\subsection{Fundamental Natural Language Processing Techniques in Comment Generation}

Software engineering tools often use part-of-speech (POS) taggers to identify the POS of a word and tag the word as a noun, verb, preposition, etc. and then chunk the tagged words into grammatical phrases to help distinguish the semantics of the component words.
Traditional taggers trained on natural language texts work well on the news and natural language artifacts, but their accuracy on source code reduces as the input moves farther away from the highly structured sentences found in traditional newswire articles.

There have been researches towards building appropriate POS taggers for software engineering researches. Binkley et al.~\cite{binkley2011improving} presented a POS tagger for field names in source code. They produce sentences with templates. Falleri et al.~\cite{falleri2010automatic} use the TreeTagger~\cite{schmid2013probabilistic}, a tagger trained on English text, to perform the POS tagging. Sridhara et al.~\cite{giri08} use well known semantic similarity techniques and perform well on English text~\cite{budanitsky2006evaluating,martin2000speech,pedersen2005maximizing}. 

Many software related words are not in the source code itself, but in the various associated software artifacts, such as the online forum, bug reports, commit logs, email communications, etc. Tian et al.~\cite{tian2014automated} developed an automatic approach that builds a software specific WordNet like database by leveraging the contents of posts in Stack Overflow. They measure the similarity of words by computing the similarities of the weighted co-occurrences of these words in the textual corpus. In addition to the work of mining semantically-similar Words, Wang et al.~\cite{wu2009learning} infer semantically related tags from FreeCode. Falleri et al.~\cite{falleri2010automatic} showed how to extract a network of identifiers connected by “is-more-general-than” or “is-a-part-of” relationships from source code. 
 
Some researchers have developed automatic techniques to capture co-occurring word pairs~\cite{maarek1991information,manning1999foundations}, but co-occurrences do not capture the information about the nature of the relationship between words beyond that the words co-occur in the same context. 
However, others have leveraged the POS tagging to analyze identifiers and build models of the usage of words in identifiers. 
Shepherd et al. showed that representing method identifiers as the verb and direct object pairs can improve search by focusing on the actions and what they action upon~\cite{shepherd07}.

\section{Challenges and Opportunities}

One of the drawbacks of the current approach is that they only generate descriptive comments. These comments are commonly used but do not reflect the knowledge beyond the source code. The comment that provides developers design intuition and thoughts are equally meaningful as description comments. There has been work about \texttt{TODO} comments, but such kinds of work are very limited. In addition, different coders may need different kinds of comments to understand the source code. Intelligent and customized comments are in high demand.


In the current state, developers often use their own dataset as their subjects in the studies. The techniques using a different dataset is often difficult to compare with another technique. Moreover, many research uses human evaluation to judge the quality of the results. This increases the difficulty to build a standard benchmark.


Previous work on comment generation in the high level action approach often use heuristics to come up with rules. One reason is the lack of high-quality training data. As common sense, developers write comments less frequently than expected. Building a high-quality dataset would not only help researchers develop an automated approach based on machine learning, but also lead to more standard matrics for evaluation and comparison.


Recurrent Neural Networks (RNNs) are gaining a lot of attention in recent years because it has shown great promise in many natural language processing tasks. Despite their popularity, to my knowledge, there has not been work using RNNs to generate comments. The nature of RNNs advanced natural language generation and translation. It seems fit in comment generation also, as descriptive comments are essentially summarization and translation of source code to natural language. 


As labeled data is always difficult to obtain. High-quality data limit the application of supervised learning methods. In machine learning, there have been many unsupervised methods that might be explored on source code data. These methods include clustering, latent semantic analysis, latent Dirichlet allocation, etc. These methods could potentially provide labels for the supervised learning methods.

Besides the progress made by many types of research in the area of comment generation, the state of the art techniques are still very limited.  This survey serves only a start of exploration of this area. 

\bibliographystyle{plain}
\bibliography{main} 

\end{document}